OpenStudy (kj4uts):

If f(x)=4x, g(x)=5x-3, and h(x)=√x, find (fogoh)(4)?

OpenStudy (kj4uts):

I attached the question!

OpenStudy (anonymous):

plug 4 into x in the g(x) equation and take that output and plug that into the f(X) equation

OpenStudy (anonymous):

(fogoh)=4(5(sqrt x)-3) =20 sqrt x-12 x=4, fogoh= 40-12=28

OpenStudy (kj4uts):

Thank you but im still a little confused :(

OpenStudy (anonymous):

Find f(g(h(x)))